OpenStudy (firechild17):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half. y − 6 = one half(x + 3) y = one halfx − 6 y + 3 = one half(x − 6) x − 2y = 12

OpenStudy (firechild17):

I dont understand this at all...pleas help me....

OpenStudy (amorfide):

standard point-slope form is y=mx+c m is the slope slope is 1/2 y=1/2x + c (6,-3) x=6, y=-3 -3=6/2 + c work out c then sub c into y=1/2x + c

OpenStudy (firechild17):

I thought the point-slope form was y=mx+b @amorfide

OpenStudy (amorfide):

a, b, c, d, e take your pick, they are just constants, in england we use c. Sorry, if you use b, just use b instead of c y=mx+b y=1/2x + b

OpenStudy (amorfide):

solve for b

OpenStudy (firechild17):

ok

OpenStudy (firechild17):

so its C?

OpenStudy (amorfide):

b=-6 y=halfx - 6 so it would be B

OpenStudy (firechild17):

thanks